Circular No. 6590

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) C. M. Wright, Max-Planck-Institut für Extraterrestrische Physik

(MPE); C. H. Smith, University College, Australian Defence Force Academy; A. Krabbe, MPE; T. Lehmann, University of Jena; and T. Böker, Space Telescope Science Institute, report: “Images in a broad-band 8–13μm filter (N band) with the MPG/ESO 2.2-m telescope (+ mid-infrared camera MANIAC) on 1996 Oct. 28.8, 30.8, 31.8, and Nov. 1.8 UT show a clear fan-shaped structure extending ∼ 30′′ in length, with a symmetry axis pointing almost due east. The size and orientation appeared stable over all four nights, and there was no obvious jet activity. N-band fluxes in a 21′′.6-diameter aperture were 14.7 ± 0.3, 16.7 ± 0.1, 18.4 ± 0.1, and 18.0 ± 0.1 × 10^{−17} W cm^{−2} μm^{−1}, respectively, on the four nights. Images in a broad–band 17.3–22.7-μm filter (Q band) on Oct. 30.8 and 31.8 and Nov. 1.8 showed a morphology consistent with that in the N band; Q-band fluxes (21′′.6 aperture) were 16.0 ± 0.2, 12.6 ± 0.3, and 14.8 ± 0.4 × 10^{−17} W cm^{−2} μm^{−1}. (Uncertainties are statistical; absolute calibration is ∼ 30 percent). N-to-Q color temperatures (with no correction for silicate emission) on Oct. 30.8, 31.8, and Nov. 1.8 were 207, 233, and 218 ± 15 K, respectively. We see little evidence for variability. On Nov. 1.8, fluxes in a 21′′.6-diameter aperture in narrow-band 8.025-, 10.53-, and 12.95-μm filters were 2.1 ± 0.5, 11.1 ± 0.4, and 7.2 ± 0.3 × 10^{−17} W cm^{−2} μm^{−1}, respectively. The 8.0–13.0-μm color temperature is 188 ± 12 K, only marginally above the 175 K expected from a blackbody in radiative equilibrium at 2.5 AU from the sun. The silicate feature, as judged from the 10.5-μm data, is a factor of 2.1 ± 0.2 above the 188 K continuum.”

A. Lovell, P. Schloerb, C. DeVries, J. Dickens, W. Irvine, and M. Senay, University of Massachusetts, write: “Maps (field size 4′ × 4′; resolution 1′) obtained during Mar. 11–13 with the FCRAO 14-m antenna (+ focalplane array receiver) of the HCN J=1–0, CS J=2–1, and HCO^{+} J=1–0 transitions show that HCN and CS emission peaks at the position of the nucleus, but the HCO^{+} emission is highly asymmetric (peaking at a point offset 1′.5 from the nucleus in p.a. ∼ 320^{o}; anti-sunward direction was ∼ p.a. 340^{o}). There is a small HCN extension toward the p.a. of peak HCO^{+} brightness. The HCN and CS lines have widths of 2.0 km/s and are centered on the velocity of the nucleus. However, the HCO^{+} line is broader at the nucleus position (width 3.2 km/s), and it is redshifted by 1.0 km/s. The HCO^{+} line width increases to 4.1 km/s at its peak-intensity position, and there is a small increase in the observed redshift of ∼ 0.2 km/s.”

1997 March 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green